Va:

A washer and a dryer cost $1000 combined. The cost of the washer is three times the cost of the dryer. What is the cost of the dryer?

Va:

anyone?

jhonyy9:

let the cos of the dryer be x and of a washer y x +y = 1000 y = 3x solve this system for x and y and will get sure easy the correct answers @supie

jhonyy9:

@Downpour17 opinion pls. ?

Downpour17:

Honestly, I think that's right. It looks accurate with the information you have.

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Downpour17 Honestly, I think that's right. It looks accurate with the information you have. \(\color{#0cbb34}{\text{End of Quote}}\) ty

CripQUEZZ:

Ok this pretty easy first get your key X = dryer 3x = washer Then set it up X + 3x = 1000 Add common variables 3x + x = 4x 4x = 1000 Then do the opposite of what 4x is doing and divide it on both sides 4x/4 = 1000/4 The (4x/4) cancels out and your left with (x = 1000/4) Then simplify 1000/4 = 250 So x = 250 So the dryer is $250 Then get the washer and since x = 250 multiply it by the washer 3(250) = 750 So the dryer is 250 and the washer is 750

CripQUEZZ:

@jhonyy9 is that right?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ @jhonyy9 is that right? \(\color{#0cbb34}{\text{End of Quote}}\) yes ofc

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ @jhonyy9 is that right? \(\color{#0cbb34}{\text{End of Quote}}\) yes ofc \(\color{#0cbb34}{\text{End of Quote}}\) Ok thx